Searching Local Periodic Families of ODEs in a Normal Form Method

Victor F. Edneral Institute for Nuclear Physics Moscow State University 119899 Vorobievy Gory, Moscow, Russia Email: edneral@theory.npi.msu.su

Resume: The report describes the method of approximating local periodic families of autonomous systems of ODE with polynomial nonlinearity in right sides by Fourier-power series. There are several examples of a treatment of systems second and forth order.